# Enhancing Topological Dependencies in Spatio-Temporal Graphs with Cycle Message Passing Blocks

### v2

@@ 1 Introduction
Spatio-temporal forecasting predicts future events or states involving both spatial and temporal components, with traffic forecasting being one of the most representative examples among these problems. Traffic data is frequently conceptualized as a spatio-temporal graph, where road connections and traffic flows gleaned from sensors are represented as edges and nodes [1, 2], and traffic forecasting aims to predict future traffic in road networks based on preceding traffic data [3, 4, 5, 6, 7]. Given this, many turn to Graph Neural Networks (GNNs), especially those using Message Passing Neural Networks (MPNNs), to study this data. GNNs based on MPNNs have been increasingly adopted to learn the complex vector representations of spatio-temporal graphs, capturing intricate dependencies in traffic datasets [8, 9, 10]. However, traditional message passing-based methodologies have exhibited limitations, including the inability to adequately account for temporal variations in traffic data and over-smoothing problems [11]. Following the success of Transformers in various domains, they have been employed either independently or synergetically with MPNNs to address these challenges in traffic forecasting [12, 13, 11], some implementations of which are coupled with preprocessing techniques like Dynamic Time Warping (DTW) [14], showcasing competitive performance outcomes [11].
@@
Building on this mathematical foundation, we propose a simple yet unique model, Cycle to Mixer (Cy2Mixer), that integrates gated Multi-Layer Perceptron (gMLP) [15] and MPNNs supplemented with topological non-trivial invariants in graphs, to enhance predictive accuracy. Our Cy2Mixer layer comprises of three key components: (a) Temporal block: Seizes the temporal characteristics of all nodes. (b) Spatial message-passing block: Encompasses spatial relationships along with local neighborhoods. (c) Cycle message-passing block: Captures supplementary information among nodes in cyclic subgraphs. Notably, our cycle message-passing block aims to encapsulate topological features, enabling a finer comprehension of the intricate connectivity patterns within the network inspired by Cy2C-GNNs [16, 17]. We validate our methodology using traffic datasets, showcasing the competitive edge of our proposed models. Furthermore, a qualitative assessment elucidates the efficacy of topological invariants in enhancing the Cy2Mixer’s predictive performance as demonstrated in Figure 1, showing improved accuracy for nodes connected through cyclic subgraphs when using the clique adjacency matrix. While our experiments focus on traffic datasets as representative spatio-temporal datasets, we also conducted experiments to demonstrate that our methodology performs well on other spatio-temporal datasets. Our work demonstrates that our approach is not limited to a single task but performs effectively across various tasks, while also evaluating the efficiency of preprocessing time.
Figure 1: Prediction example of Cy2Mixer in the PEMS04 dataset. While Cy2Mixer shows similar performance between Node 170 and Node 173, which are connected via the adjacency matrix $A$, Cy2Mixer exhibits superior performance at Node 0 when utilizing the clique adjacency matrix $A_{C}$, indicating the effectiveness of the cycle block in capturing cyclic subgraph relationships.
@@ 2 Related Works
In recent years, deep learning approaches have emerged as powerful tools in the spatio-temporal modeling by leveraging their ability to automatically learn complex, nonlinear spatio-temporal patterns, and traffic forecasting is one of the primary applications of spatio-temporal modeling. Convolutional Neural Networks (CNNs) have been effectively applied to traffic forecasting to capture spatio-temporal patterns [18, 3, 19]. With the success of Transformers in fields like natural language processing and computer vision, owing to their ability to model complex relationships over long sequences without relying on recurrent structures, they have been effectively adapted to traffic forecasting by leveraging self-attention mechanisms to capture long-range spatio-temporal dependencies [13, 11, 20]. Meanwhile, to capture the dynamic spatial correlations within traffic networks, Graph Neural Networks (GNNs) have become a prominent approach for traffic forecasting due to their ability to model complex dependencies [8, 21, 22, 10], and some studies have integrated self-attention mechanisms with GNNs to further enhance the modeling of spatial relationships within traffic networks [12, 23]. Some recent studies adapt traffic graph structures through learnable adjacency matrices and probabilistic modeling [24, 25] and integrate hierarchical topological views to capture multi-level spatial patterns [26, 27]. However, many existing approaches that incorporate spatial information in traffic networks have relied on either limited methods or overly heuristic strategies with auxilary information from road networks, which may be challenging to apply to other spatio-temporal tasks. These methods often struggle to comprehensively capture the full complexity of spatial dependencies, leading to suboptimal modeling of traffic dynamics.
@@ 3.1 Message passing neural networks (MPNNs)
where $\text{where }h_{u}^{\prime}=\text{AGGREGATE}_{v}^{(m)}\left(\left\{\!\!\left\{h_{u}^{(m-1)}\;|\;u\in N(v)\right\}\!\!\right\}\right).$ (2)
@@
gMLP has been shown to achieve comparable performance to Transformer models across diverse domains, including computer vision and natural language processing, with improved efficiency [15]. With the given input $X\in\mathbb{R}^{n\times d}$, where $n$ denotes sequence length and $d$ denotes dimension, it can be defined as:
@@
where $\sigma$ represents the activation function, which is GeLU [28] in this context, and $U$ and $V$ correspond to linear projections based on the channel (feature) dimensions, serving roles similar to the feed-forward network in Transformer. The layer $s(\cdot)$ corresponds to the Spatial Gating Unit, which is responsible for capturing cross-token interactions. We construct the aforementioned layer as:
@@ 3.3 Traffic Prediction Problem
Traffic network can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},A,A_{C})$, where $\mathcal{V}=\{v_{1},\cdots,v_{N}\}$ denotes the set of $N$ nodes representing sensors within the traffic network ($|V|=N$). Next, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, and the adjacency matrix $A$ of the network $\mathcal{G}$ can be obtained based on the distances between nodes. Additionally, $A_{C}$ is the clique adjacency matrix of the network, which is incorporated into the GNN architecture. The detailed procedure for obtaining the clique adjacency matrix is provided in Appendix A.7. The matrices $A\in\mathbb{R}^{N\times N}$ and $A_{C}\in\mathbb{R}^{N\times N}$ are time-independent input variables since the structure remains unchanged over time. As will be noted in Section 4 and Appendix A.6, cyclic substructures of traffic network are crucial in understanding additional conditions imposed in accurately predicting traffic signals. As demonstrated in previous studies [16], utilizing clique adjacency matrices $A_{C}$ can effectively boost capabilities of conventional deep learning architectures in incorporating cyclic substructures of traffic networks. As such, we use the clique adjacency matrix $A_{C}$ as additional input to our traffic network data.
@@ 4 Mathematical Backgrounds
One of the prominent ways to mathematically quantify the effectiveness and discerning capabilities of GNNs is to interpret the graph dataset as a collection of 1-dimensional topological spaces $\mathcal{G}:=\{G_{i}\}_{i}$, and identify the given GNNs as a function $GNN:\mathcal{G}\to\mathbb{R}^{k}$ which represents graphs in a given dataset as real $k$-dimensional vectors. These approaches were carefully executed in previous literature, which focused on pinpointing the discernability of various architectural designs for improving GNNs, as seen in [16, 29, 30, 31, 32]. Throughout these references, the characterizations of topological invariants of such 1-dimensional topological spaces provided grounds for verifying whether such GNNs can effectively capture geometric non-trivial properties of graph datasets, such as cyclic substructures or connected components of graphs. In light of these previous studies, it is hence of paramount interest to reinterpret temporal graph datasets as a collection of higher dimensional topological spaces and understand what non-trivial properties of such topological spaces the novel deep learning techniques processing temporal graph datasets should encapsulate.
@@
One of the topological invariants that conventional GNNs fail to incorporate effectively is the cyclic substructures of graphs [33, 29]. This is because GNNs can only distinguish any pairs of two graphs $G$ and $H$ up to isomorphism of their universal covers or unfolding trees. To overcome this issue, one can substitute cyclic subgraphs with cliques. An approach inspired by the theory of covering spaces, the strategy alters the geometry of universal covers to allow GNNs to encapsulate cyclic structures effectively [16]. We implement the application of the analogous operation to temporal traffic data by utilizing the clique adjacency matrix $A_{C}$.The clique adjacency matrix $A_{C}$ allows room for GNNs to determine whether additional interactions among nodes of $\mathcal{G}$, whose measurements are taken at varying moments of time, are relevant components for forecasting traffic signals. This allows us to capture potential effects of topological invariants of traffic networks on temporal traffic measurements, as suggested from Theorem 4.1. Figure 2 illustrates how lifting a cyclic subgraph of $\mathcal{G}$ to a new cyclic subgraph of $\mathcal{G}\times I$ of varying temporal instances can enrich our prediction of future traffic signals. The two cycles are subgraphs colored in red. The temporal cyclic subgraph spans over time $t=t_{0}$ and the end time $t=t_{e}$ of the dataset. By changing the colored cyclic subgraph into cliques, we add edges to nodes on the cyclic subgraph lying in two different instances of time. These edges establish relations among measurements taken at different nodes and time periods.
@@ 5.1 Embedding Layer
Our model follows the structure of the embedding layer from Liu et al. [20]. Given the time-series data $X_{t-T+1:t}$, the feature embedding is defined as $X_{\textit{feat}}=\text{FC}(X_{t-T+1:t})$, where FC represents a fully connected layer and $X_{\textit{feat}}\in\mathbb{R}^{T\times N\times d_{f}}$. Additionally, to account for both weekly and daily periodicities, the temporal embedding $X_{\textit{temp}}\in\mathbb{R}^{T\times N\times d_{t}}$ is produced. This is achieved by referencing the day-of-week and timestamp-of-day data to extract two distinct embeddings, which are subsequently concatenated and broadcasted to generate the final temporal embedding with dimension $d_{t}$. Furthermore, to address diverse temporal patterns specific to each node, a learnable parameter is introduced in the form of an adaptive embedding of dimension $d_{a}$, represented as $X_{\textit{ast}}\in{\mathbb{R}^{T\times N\times d_{a}}}$, to capture and reflect latent, unobserved information within the temporal and spatial patterns of each node. These three embeddings are concatenated, forming the output $H\in\mathbb{R}^{T\times N\times d_{h}}$ of the embedding layer, where the dimension $d_{h}$ is equal to $d_{f}+d_{t}+d_{a}$.
$H=X_{\textit{feat}}||X_{\textit{temp}}||X_{\textit{ast}}.$ (8)
@@ 5.2 Cy2Mixer Encoder Layer
These blocks employ the MPNN for their projection function in the Gating Unit. Notably, the spatial message-passing block uses the standard adjacency matrix, $A$, whereas the cycle message-passing block operates with the clique adjacency matrix, $A_{C}$. The clique adjacency matrix, proposed by  [16], represents the bases of cycles (or the first homological invariants of the graph [34]) in a suitable form that enables GNNs to effectively process the desired topological features. Following a process similar to the Temporal block, the $\tilde{Z}$ values for both blocks are expressed as:
@@
where $\tilde{Z}_{\textit{spatial}}\in\mathbb{R}^{T\times N\times d_{h}}$ and $\tilde{Z}_{\textit{cycle}}\in\mathbb{R}^{T\times N\times d_{h}}$. The final output of each block, denoted as $Y$, is calculated by applying a linear projection to the output of Gating Unit $\tilde{Z}$, similar to the process in Section 2.2. Considering the effective performance outcomes of incorporating tiny attention into each block in prior research [15], we have adopted the same structure in our model. Then we concatenate all of three outputs, namely, $Y_{\textit{temporal}},Y_{\textit{cycle}}$, and $Y_{\textit{spatial}}$, using a feed-forward network as:
$Y_{\textit{out}}=\text{FC}\left(Y_{\textit{temporal}}||Y_{\textit{cycle}}||Y_{\textit{spatial}}\right),\quad Y_{\textit{out}}\in\mathbb{R}^{T\times N\times d_{h}}.$ (12)
@@ 5.3 Output Layer
Table 1: Traffic flow prediction results for PEMS04, PEMS07, PEMS08, NYTaxi, TDrive, and CHBike. All prediction results in gray are cited from available results obtained from pre-existing publications [20]. Highlighted are the top first and second results.
@@ 6 Experiments
In this study, we evaluate our proposed approach against various established baseline methods in traffic forecasting. We include GNN-based methods such as GWNet [22] and STGNCDE [10]. We also include GMAN [12], ASTGNN [23], PDFormer [11] and STAEFormer [20], all of which are self-attention-based models designed for the same task as ours.
@@
Cy2Mixer (PEMS04) MAE RMSE MAPE Temporal Block Spatial Block Cycle Block $A$ $A_{C}$ Tiny attention 18.55 30.31 12.20% 22.53 36.33 15.36% 18.81 30.65 12.86% 18.85 30.68 12.72% 18.56 30.33 12.32% 18.29 30.05 11.95% 18.36 30.15 12.05% 18.14 30.02 11.93%

### v3

@@ 1 Introduction
Spatio-temporal forecasting predicts future events or states involving both spatial and temporal components, with traffic forecasting being one of the most representative examples among these problems. Traffic data is frequently conceptualized as a spatio-temporal graph, where road connections and traffic flows gleaned from sensors are represented as edges and nodes [diao2019dynamic, wang2020traffic], and traffic forecasting aims to predict future traffic in road networks based on preceding traffic data [li2017diffusion, deng2021st, shao2022spatial, shao2022decoupled, zhu2023correlation]. Given this, many turn to Graph Neural Networks (GNNs), especially those using Message Passing Neural Networks (MPNNs), to study this data. GNNs based on MPNNs have been increasingly adopted to learn the complex vector representations of spatio-temporal graphs, capturing intricate dependencies in traffic datasets [yu2017spatio, li2021spatialtemporal, choi2022graph]. However, traditional message passing-based methodologies have exhibited limitations, including the inability to adequately account for temporal variations in traffic data and over-smoothing problems [jiang2023pdformer]. Following the success of Transformers in various domains, they have been employed either independently or synergetically with MPNNs to address these challenges in traffic forecasting [zheng2020gman, xu2020spatial, jiang2023pdformer], some implementations of which are coupled with preprocessing techniques like Dynamic Time Warping (DTW) [berndt1994using], showcasing competitive performance outcomes [jiang2023pdformer].
@@
Building on this mathematical foundation, we propose a simple yet unique model, Cycle to Mixer (Cy2Mixer), that integrates gated Multi-Layer Perceptron (gMLP) [liu2021pay] and MPNNs supplemented with topological non-trivial invariants in graphs, to enhance predictive accuracy. Our Cy2Mixer layer comprises of three key components: (a) Temporal block: Seizes the temporal characteristics of all nodes. (b) Spatial message-passing block: Encompasses spatial relationships along with local neighborhoods. (c) Cycle message-passing block: Captures supplementary information among nodes in cyclic subgraphs. Notably, our cycle message-passing block aims to encapsulate topological features, enabling a finer comprehension of the intricate connectivity patterns within the network inspired by Cy2C-GNNs [choi2022cycle, choi2024topology]. We validate our methodology using traffic datasets, showcasing the competitive edge of our proposed models. Furthermore, a qualitative assessment elucidates the efficacy of topological invariants in enhancing the Cy2Mixer’s predictive performance as demonstrated in Figure 1, showing improved accuracy for nodes connected through cyclic subgraphs when using the clique adjacency matrix. While our experiments focus on traffic datasets as representative spatio-temporal datasets, we also conducted experiments to demonstrate that our methodology performs well on other spatio-temporal datasets. Our work demonstrates that our approach is not limited to a single task but performs effectively across various tasks, while also evaluating the efficiency of preprocessing time.
Figure 1: Prediction example of Cy2Mixer in the PEMS04 dataset. While Cy2Mixer shows similar performance between Node 170 and Node 173, which are connected via the adjacency matrix AA, Cy2Mixer exhibits superior performance at Node 0 when utilizing the clique adjacency matrix $A_{C}$, indicating the effectiveness of the cycle block in capturing cyclic subgraph relationships.
@@ 2 Related Works
In recent years, deep learning approaches have emerged as powerful tools in the spatio-temporal modeling by leveraging their ability to automatically learn complex, nonlinear spatio-temporal patterns, and traffic forecasting is one of the primary applications of spatio-temporal modeling. Convolutional Neural Networks (CNNs) have been effectively applied to traffic forecasting to capture spatio-temporal patterns [zhang2017deep, li2017diffusion, lin2020preserving]. With the success of Transformers in fields like natural language processing and computer vision, owing to their ability to model complex relationships over long sequences without relying on recurrent structures, they have been effectively adapted to traffic forecasting by leveraging self-attention mechanisms to capture long-range spatio-temporal dependencies [xu2020spatial, jiang2023pdformer, liu2023spatio]. Meanwhile, to capture the dynamic spatial correlations within traffic networks, Graph Neural Networks (GNNs) have become a prominent approach for traffic forecasting due to their ability to model complex dependencies [yu2017spatio, pan2019urban, wu2019graph, choi2022graph], and some studies have integrated self-attention mechanisms with GNNs to further enhance the modeling of spatial relationships within traffic networks [zheng2020gman, guo2021learning]. Some recent studies adapt traffic graph structures through learnable adjacency matrices and probabilistic modeling [wu2020connecting, shang2021discrete] and integrate hierarchical topological views to capture multi-level spatial patterns [wu2020learning, ma2023rethinking]. However, many existing approaches that incorporate spatial information in traffic networks have relied on either limited methods or overly heuristic strategies with auxilary information from road networks, which may be challenging to apply to other spatio-temporal tasks. These methods often struggle to comprehensively capture the full complexity of spatial dependencies, leading to suboptimal modeling of traffic dynamics.
@@ 3.1 Message passing neural networks (MPNNs)
$\text{where }h_{u}^{\prime}=\text{AGGREGATE}_{v}^{(m)}\left(\left\{\!\!\left\{h_{u}^{(m-1)}\;|\;u\in N(v)\right\}\!\!\right\}\right).$ (2)
@@
gMLP has been shown to achieve comparable performance to Transformer models across diverse domains, including computer vision and natural language processing, with improved efficiency [liu2021pay]. With the given input $X\in\mathbb{R}^{n\times d}$, where $n$ denotes sequence length and $d$ denotes dimension, it can be defined as:
@@
where $\sigma$ represents the activation function, which is GeLU [hendrycks2023gaussian] in this context, and $U$ and $V$ correspond to linear projections based on the channel (feature) dimensions, serving roles similar to the feed-forward network in Transformer. The layer $s(\cdot)$ corresponds to the Spatial Gating Unit, which is responsible for capturing cross-token interactions. We construct the aforementioned layer as:
@@ 3.3 Traffic Prediction Problem
Traffic network can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},A,A_{C})$, where $\mathcal{V}=\{v_{1},\cdots,v_{N}\}$ denotes the set of $N$ nodes representing sensors within the traffic network ($|V|=N$). Next, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, and the adjacency matrix $A$ of the network $\mathcal{G}$ can be obtained based on the distances between nodes. Additionally, $A_{C}$ is the clique adjacency matrix of the network, which is incorporated into the GNN architecture. The detailed procedure for obtaining the clique adjacency matrix is provided in Appendix A.7. The matrices $A\in\mathbb{R}^{N\times N}$ and $A_{C}\in\mathbb{R}^{N\times N}$ are time-independent input variables since the structure remains unchanged over time. As will be noted in Section 4 and Appendix A.6, cyclic substructures of traffic network are crucial in understanding additional conditions imposed in accurately predicting traffic signals. As demonstrated in previous studies [choi2022cycle], utilizing clique adjacency matrices $A_{C}$ can effectively boost capabilities of conventional deep learning architectures in incorporating cyclic substructures of traffic networks. As such, we use the clique adjacency matrix $A_{C}$ as additional input to our traffic network data.
@@ 4 Mathematical Backgrounds
One of the prominent ways to mathematically quantify the effectiveness and discerning capabilities of GNNs is to interpret the graph dataset as a collection of 1-dimensional topological spaces $\mathcal{G}:=\{G_{i}\}_{i}$, and identify the given GNNs as a function $GNN:\mathcal{G}\to\mathbb{R}^{k}$ which represents graphs in a given dataset as real $k$-dimensional vectors. These approaches were carefully executed in previous literature, which focused on pinpointing the discernability of various architectural designs for improving GNNs, as seen in [choi2022cycle, bodnar21cw, horn2021topological, park2022pwlr, bernardez2024icml]. Throughout these references, the characterizations of topological invariants of such 1-dimensional topological spaces provided grounds for verifying whether such GNNs can effectively capture geometric non-trivial properties of graph datasets, such as cyclic substructures or connected components of graphs. In light of these previous studies, it is hence of paramount interest to reinterpret temporal graph datasets as a collection of higher dimensional topological spaces and understand what non-trivial properties of such topological spaces the novel deep learning techniques processing temporal graph datasets should encapsulate.
@@
One of the topological invariants that conventional GNNs fail to incorporate effectively is the cyclic substructures of graphs [xu2019powerful, bodnar21cw]. This is because GNNs can only distinguish any pairs of two graphs $G$ and $H$ up to isomorphism of their universal covers or unfolding trees. To overcome this issue, one can substitute cyclic subgraphs with cliques. An approach inspired by the theory of covering spaces, the strategy alters the geometry of universal covers to allow GNNs to encapsulate cyclic structures effectively [choi2022cycle]. We implement the application of the analogous operation to temporal traffic data by utilizing the clique adjacency matrix $A_{C}$.The clique adjacency matrix $A_{C}$ allows room for GNNs to determine whether additional interactions among nodes of $\mathcal{G}$, whose measurements are taken at varying moments of time, are relevant components for forecasting traffic signals. This allows us to capture potential effects of topological invariants of traffic networks on temporal traffic measurements, as suggested from Theorem 4.1. Figure 2 illustrates how lifting a cyclic subgraph of $\mathcal{G}$ to a new cyclic subgraph of $\mathcal{G}\times I$ of varying temporal instances can enrich our prediction of future traffic signals. The two cycles are subgraphs colored in red. The temporal cyclic subgraph spans over time $t=t_{0}$ and the end time $t=t_{e}$ of the dataset. By changing the colored cyclic subgraph into cliques, we add edges to nodes on the cyclic subgraph lying in two different instances of time. These edges establish relations among measurements taken at different nodes and time periods.
@@ 5.1 Embedding Layer
Our model follows the structure of the embedding layer from liu2023spatio. Given the time-series data $X_{t-T+1:t}$, the feature embedding is defined as $X_{\textit{feat}}=\text{FC}(X_{t-T+1:t})$, where FC represents a fully connected layer and $X_{\textit{feat}}\in\mathbb{R}^{T\times N\times d_{f}}$. Additionally, to account for both weekly and daily periodicities, the temporal embedding $X_{\textit{temp}}\in\mathbb{R}^{T\times N\times d_{t}}$ is produced. This is achieved by referencing the day-of-week and timestamp-of-day data to extract two distinct embeddings, which are subsequently concatenated and broadcasted to generate the final temporal embedding with dimension $d_{t}$. Furthermore, to address diverse temporal patterns specific to each node, a learnable parameter is introduced in the form of an adaptive embedding of dimension $d_{a}$, represented as $X_{\textit{ast}}\in{\mathbb{R}^{T\times N\times d_{a}}}$, to capture and reflect latent, unobserved information within the temporal and spatial patterns of each node. These three embeddings are concatenated, forming the output $H\in\mathbb{R}^{T\times N\times d_{h}}$ of the embedding layer, where the dimension $d_{h}$ is equal to $d_{f}+d_{t}+d_{a}$.
[…]
@@ 5.2 Cy2Mixer Encoder Layer
These blocks employ the MPNN for their projection function in the Gating Unit. Notably, the spatial message-passing block uses the standard adjacency matrix, $A$, whereas the cycle message-passing block operates with the clique adjacency matrix, $A_{C}$. The clique adjacency matrix, proposed by  [choi2022cycle], represents the bases of cycles (or the first homological invariants of the graph [paton1969algorithm]) in a suitable form that enables GNNs to effectively process the desired topological features. Following a process similar to the Temporal block, the $\tilde{Z}$ values for both blocks are expressed as:
@@
where $\tilde{Z}_{\textit{spatial}}\in\mathbb{R}^{T\times N\times d_{h}}$ and $\tilde{Z}_{\textit{cycle}}\in\mathbb{R}^{T\times N\times d_{h}}$. The final output of each block, denoted as $Y$, is calculated by applying a linear projection to the output of Gating Unit $\tilde{Z}$, similar to the process in Section 2.2. Considering the effective performance outcomes of incorporating tiny attention into each block in prior research [liu2021pay], we have adopted the same structure in our model. Then we concatenate all of three outputs, namely, $Y_{\textit{temporal}},Y_{\textit{cycle}}$, and $Y_{\textit{spatial}}$, using a feed-forward network as:
[…]
@@ 5.3 Output Layer
Table 1: Traffic flow prediction results for PEMS04, PEMS07, PEMS08, NYTaxi, TDrive, and CHBike. All prediction results in gray are cited from available results obtained from pre-existing publications [liu2023spatio]. Highlighted are the top first and second results.
@@ 6 Experiments
In this study, we evaluate our proposed approach against various established baseline methods in traffic forecasting. We include GNN-based methods such as GWNet [wu2019graph] and STGNCDE [choi2022graph]. We also include GMAN [zheng2020gman], ASTGNN [guo2021learning], PDFormer [jiang2023pdformer] and STAEFormer [liu2023spatio], all of which are self-attention-based models designed for the same task as ours.
@@
Cy2Mixer (PEMS04) MAE RMSE MAPE Temporal Block Spatial Block Cycle Block AA $A_{C}$ Tiny attention 18.55 30.31 12.20% 22.53 36.33 15.36% 18.81 30.65 12.86% 18.85 30.68 12.72% 18.56 30.33 12.32% 18.29 30.05 11.95% 18.36 30.15 12.05% 18.14 30.02 11.93%